AM3 -Esercitazione 3 A.A. 2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 8 marzo 2004

Esercizio 1

Risolvere le seguenti equazioni differenziali ordinarie a coefficienti costanti:

1. ü − 3u̇ + 2u = 0;

2. ü − 2u̇ + u = 0;

3. u^{(3)} − 2u^{(2)} + 2u^{(1)} = 0;

4. ü + ω^{2}u = 0 ω ∈ ℝ;

5. u̇ − u = cos t.